# Patterns Over Principles: The Fragility of Inductive Reasoning in LLMs under Noisy Observations

### v1

@@ Abstract
Inductive reasoning, a cornerstone of human cognition, enables generalization from limited data but hasn’t yet been fully achieved by large language models (LLMs). While modern LLMs excel at reasoning tasks, their ability to maintain stable and consistent rule abstraction under imperfect observations remains underexplored. To fill this gap, in this work, we introduce Robust Rule Induction, a task that evaluates LLMs’ capability in inferring rules from data that are fused with noisy examples. To address this task, we further propose Sample-steered Rule Refinement (SRR), a method enhancing reasoning stability via observation diversification and execution-guided feedback. Experiments across arithmetic, cryptography, and list functions reveal: (1) SRR outperforms other methods with minimal performance degradation under noise; (2) Despite slight accuracy variation, LLMs exhibit instability under noise (e.g., accuracy change with only consistent score); (3) Counterfactual task gaps highlight LLMs’ reliance on memorized patterns over genuine abstraction. Our findings challenge LLMs’ reasoning robustness, revealing susceptibility to hypothesis drift and pattern overfitting, while providing empirical evidence critical for developing human-like inductive systems. Code and data are available at https://github.com/lcy2723/Robust-Rule-Induction.
@@ 1 Introduction
Despite increasing attention to LLM inductive reasoning capabilities, current works exhibit limitations that obscure true inductive reasoning robustness. First, in prevailing evaluations Alet et al. (2021); Gendron et al. (2024), models just predict outputs for novel inputs given exemplars, bypassing explicit rule verification. It fails to diagnose where and why rule induction fails. Existing studies that do generate intermediate rules either disregard noisy learning conditions or assess performance solely through aggregate metrics like task accuracy gap. This neglects instance-level reasoning consistency—whether models maintain stable rule interpretations when exposed to conflicting patterns, which is a key indicator of human-like robust induction. Besides, prior methods often repurpose existing benchmarks (e.g., SCAN; Lake and Baroni 2018, ARC; Chollet 2019) without modification, where potential data contamination may undermine validity as models could use memorized solutions.
To address these gaps, we introduce a novel task of robust rule induction, which challenges models to identify underlying functions from input-output exemplars containing controlled noise injections. Our approach features three key innovations: (1) An evaluation pipeline from data synthesis to automatic evaluation, requiring models to explicitly output rules and enabling direct validation through programmatic execution. (2) A multi-dimensional robustness assessment with conventional accuracy and instance-level metrics. (3) A sample-steered refinement (SRR) methodology that improves both the capability and stability of LLMs’ inductive reasoning through diverse sampling and diagnostic refinement.
@@ 2.1 Problem Definitions
We formulate robust rule induction as the problem of identifying latent mapping functions from input-output pairs while maintaining consistent reasoning performance under noisy conditions. While prior works Qiu et al. (2024); Wang et al. (2024) focus on basic rule discovery capabilities, we specifically investigate the robustness of inductive reasoning when the observations contain conflicting patterns. Some examples are shown in Figure 1. The observations may contain both normal and noisy examples, and the model is required to infer the correct rule regardless of the noise.
@@ 3 Sample-steered Rule Refinement
To address noisy rule induction, we propose Sample-steered Rule Refinement (SRR), a novel framework that combines hypothesis generation, contrastive evidence sampling, and iterative self-correction. This method conducts inductive reasoning through three-phase optimization: (1) contrastive hypothesis generation to bootstrap diverse rule candidates, (2) diagnostic feedback construction through evidence-aware sampling, and (3) iterative rule refinement using failure-driven corrections. The full algorithm is detailed in Algorithm 1.
@@
We first generate hypotheses from random subsets of the seen examples plus one hypothesis using the full observations. This step ensures coverage of both local patterns and global consistency. The number of examples in the subset is less than the full seen set, thus the noise in the subset has a greater impact. The sampled subset may also contain completely noise-free examples, which can better distinguish the rules.
@@
After generating the initial hypotheses, the induced rule is executed on the seen examples via an external Python interpreter. We then collect the correct and incorrect cases and sample these cases as feedback. We focus more on the incorrect cases, as they provide more information about the rule refinement, while the correct cases are used as positive feedback to reinforce the rule. The hypothesis with the highest accuracy on the seen examples is selected as the initial selected rule.
@@
At each iteration, the model receives the current selected rule with formatted feedback to generate a new rule. The new hypothesis is compared to the previous one, and the more accurate one is selected. This iterative process continues until the accuracy of the rule on the seen examples exceeds a predefined threshold or the maximum number of iterations is reached.
@@ 4.2 Datasets
We use three datasets with their subsets: Arithmetic, Cryptography, and List Functions. These datasets offer different rule induction challenges, including mathematical calculations, symbol representation, and list operations. Comprehensive statistics are shown in Table 1.
Dataset Subset # Tasks # Normal # Noise # Test Arith. 7-base 8-base 9-base Crypto. Caesar Atbash Keyboard List Func. N/A
@@
The arithmetic task, proposed by Wu et al. (2024), is a counterfactual two-digit addition task. Instead of the common base-10 system, the task uses base-8, 9, and 11 as the counterfactual setup to control the difficulty and avoid the impact of memorization over reasoning. In this task, we focus on the base-7, 8, 9 systems, and the noisy example is the common 10-based equations. The input of each example is two two-digit numbers in the corresponding base, connected by a plus sign, and the output is the sum of the two numbers. To ensure the model can reason correctly, we guarantee that there must be a carry-over in the addition process so that the model must reason about the carry-over instead of simply adding the numbers.
@@
We use three types of substitution ciphers: Caesar, Atbash, and Keyboard. The Caesar cipher is a cipher that shifts the alphabet by a fixed number of positions. The Atbash cipher is a cipher that replaces each letter with the letter symmetrically opposite in the alphabet (e.g. AZ). The Keyboard cipher replaces letters according to their positions in the alphabet with the corresponding positions on the keyboard (e.g. AQ, W). The input of each example is a word, and the output is the encrypted text. We randomly replace the characters in the output with other characters to generate noisy examples.
@@
The list functions dataset Rule (2020) evaluates the concept learning ability in the domain of cognitive science. The task is to induce a function that maps a list of numbers to another list of numbers. The dataset has tasks. Each item in the dataset corresponds to a list manipulation operation, such as sorting, reversing, or filtering. We randomly generate the examples of each task under controlled conditions to ensure the rule can be induced. The noisy examples are generated by randomly replacing the numbers in the output list with other numbers.
@@ 4.3 Robustness Under Different Noise Levels
First, contrary to conventional assumptions, noise introduction does not universally degrade performance, instead, the models exhibit performance fluctuations, demonstrating their inherent sensitivity to conflicting patterns. For example, GPT-4o achieves improved accuracy on Caesar cipher tasks at noise ( absolute improvement over clean data), while List Functions exhibit monotonic performance decay with increasing noise levels. In some cases, moderate noise improves performance, consistent with the findings of Zhou et al. (2024).
@@
Model Method Arithmetic Cryptography List Functions 7-base 8-base 9-base Caesar Atbash Keyboard GPT-4o DO CoT SC SR Ours Deepseek-V3 DO CoT SC SR Ours
Dataset GPT-4o DeepSeek-V3 SC SRR-0 SC SRR-0 Arithmetic Arithmetic Arithmetic Crypto Crypto Crypto List Functions
@@ 4.4 Method-wise Effectiveness Comparison
Direct Output may limit the model’s reasoning ability, as the model must output the rule directly without any intermediate steps. To systematically assess the robustness of different reasoning paradigms, we compare our Sample-steered Rule Refinement (SRR) method with three reasoning methods: (1) Chain of Thought (CoT; Wei et al. 2022), which decomposes reasoning into step-by-step rationales; (2) Self-Consistency (SC; Wang et al. 2023), which aggregates multi CoT trajectories through majority voting; and (3) Self-Refine (SR; Madaan et al. 2023), which iteratively improves hypotheses using self-generated feedback. Table 3 presents the task accuracy under noise and the deviation from clean data for GPT-4o and DeepSeek-V3.
@@
As shown in Table 3, our SRR framework achieves state-of-the-art performance across 13/14 task-model combinations while exhibiting minimal performance degradation ( average drop vs. - for baselines). This dual advantage stems from two mechanisms: (1) Diversity-aware hypothesis generation explores broader solution spaces through subset sampling, outperforming SC’s majority voting that amplifies the similar patterns (Table 4); (2) Execution-guided feedback leverages Python interpreters for objective error detection, circumventing LLMs’ inherent deductive limitations Chen et al. (2023b); Cheng et al. (2024).
@@
While task accuracy provides comprehensive insights, consistency scores uncover fundamental reasoning fragility. As depicted in Figure 4, the Atbash cipher task exhibits particularly low consistency despite modest accuracy changes. The slight variation in accuracy may merely be an illusion created by the combined effects of noise interference (solvedunsolved cases) and incorrect generalization (unsolvedsolved cases). Prior and concurrent work’s singular focus on accuracy fluctuations Qiu et al. (2024); Zhou et al. (2024); Mirzadeh et al. (2024); Huang et al. (2025a) overlooks this critical duality in reasoning robustness.
@@
We observe dramatic performance gaps on counterfactual tasks like base-7 and base-9 arithmetic ( vs. for DeepSeek-V3 vs. GPT-4o) and Keyboard ciphers. This observation is consistent with the findings reported in Wu et al. (2024). Diagnostic analysis of responses from LLMs shows models default to familiar templates rather than true induction. GPT-4o persistently misinterprets base-7 and base-9 addition as “base-8” or “decimal sum with constant”, while Atbash and Keyboard ciphers get erroneously classified as Caesar shifts in thinking process or self-generated feedback. Although Atbash and Caesar cipher exhibit similar levels of complexity in transformation, their performance differences are significant. These failures reveal that there is a pattern overfitting in the reasoning process. Current models’ inductive reasoning essentially operates through pattern matching rather than abstract induction. When the scarcity of counterfactual tasks in pretraining data forces models to rely on genuine rule induction rather than recitation, model performance plummets dramatically.
@@ 4.5 Extended Explorations
Comparion with DeepSeek-R1
We focus on the Atbash cipher and List Functions. Atbash provides unified task semantics yet challenges models with unfamiliar transformation logic, while List Functions captures diverse rule abstraction scenarios. Table 5 compares task accuracy and consistency scores under clean and noise conditions with DeepSeek-V3 (SRR). Notably, DeepSeek-R1 achieves higher task accuracy on both tasks. However, its consistency scores are not competitive enough, indicating unresolved instability. Detailed breakdowns of consistency scores are shown in 6. For Atbash, the comparable number of right-to-wrong and wrong-to-right suggests instability and randomness in its reasoning process. Manual inspection shows DeepSeek-R1 still interprets Atbash cipher as character shifts in failed cases, mirroring previous pattern-overfitting behavior. For List Functions, the elevated right-to-wrong rate indicates its sensitivity to input perturbations.
Dataset Model Clean Acc Noise Acc Consistency Atbash SRR DeepSeek-R1 List Func. SRR DeepSeek-R1
Dataset Model # BothR # BothW # RtoW # WtoR Atbash SRR DeepSeek-R1 List Func. SRR DeepSeek-R1
@@
Rule (2020) reports human performance on List Functions. We divide the tasks into three difficulty levels in a ratio based on the sorted mean human performance. To contrast with human-like thinking patterns, we analyze consistency across 12 trials (4 noise levels 3 runs) in Section 4.3. Figure 5 visualizes the distribution of consistency and performance per task in List Functions, stratified by difficulty. LLMs show macro-level alignment yet micro-level divergence compared to human reasoning. While LLMs broadly mirror human-like stability, showing higher consistency on both simple and hard tasks, with moderate instability on medium-difficulty tasks, their internal patterns reveal critical deviations. For hard tasks, models display unpredictability (e.g., sporadic success) rather than systematic incapacity. Simple tasks, despite their low complexity, exhibit in consistent performance unrelated to intrinsic difficulty. Notably, GPT-4o and DeepSeek-V3 demonstrate similar behavioral trends, suggesting a shared inductive bias with humans.
@@ 5 Discussion
LLMs heavily rely on prior knowledge rather than pure induction. They blend memorized pattern matching with shallow reasoning. When observations align with priors (e.g., Caesar cipher), they demonstrate pseudo-robustness. However, conflicting patterns (noise) and counterfactual scenarios (unseen rules) expose this fragility—models either default to familiar templates or enter unstable hypothesis oscillations. Even the state-of-the-art reasoning model exhibits unstable performance in unified tasks. This contrasts sharply with human few-shot learning, where true rule abstraction enables stable generalization Lake et al. (2015). Achieving human-level robust reasoning and few-shot induction requires disentangling knowledge recitation from inductive rule formation.

### v2

@@ Abstract
Inductive reasoning, a cornerstone of human cognition, enables generalization from limited data but hasn’t yet been fully achieved by large language models (LLMs). While modern LLMs excel at reasoning tasks, their ability to maintain stable and consistent rule abstraction under imperfect observations remains underexplored. To fill this gap, in this work, we introduce Robust Rule Induction, a task that evaluates LLMs’ capability in inferring rules from data that are fused with noisy examples. To address this task, we further propose Sample-steered Rule Refinement (SRR), a method enhancing reasoning stability via observation diversification and execution-guided feedback. Experiments across arithmetic, cryptography, and list functions reveal: (1) SRR outperforms other methods with minimal performance degradation under noise; (2) Despite slight accuracy variation, LLMs exhibit instability under noise (e.g., accuracy change with only consistent score); (3) Counterfactual task gaps highlight LLMs’ reliance on memorized patterns over genuine abstraction. Our findings challenge LLMs’ reasoning robustness, revealing susceptibility to hypothesis drift and pattern overfitting, while providing empirical evidence critical for developing human-like inductive systems. Code and data are available at https://github.com/HKUST-KnowComp/Robust-Rule-Induction.
@@ 1 Introduction
Despite increasing attention to LLM inductive reasoning capabilities, current works exhibit limitations that obscure true inductive reasoning robustness. First, in prevailing and concurrent evaluations Alet et al. (2021); Gendron et al. (2024); Zheng et al. (2025b), models just predict outputs for novel inputs given exemplars, bypassing explicit rule verification. It fails to diagnose where and why rule induction fails. Existing studies that do generate intermediate rules either disregard noisy learning conditions Wang et al. (2024); Liu et al. (2024) or assess performance solely through aggregate metrics like task accuracy gap Qiu et al. (2024). This neglects instance-level reasoning consistency—whether models maintain stable rule interpretations when exposed to conflicting patterns, which is a key indicator of human-like robust induction. Besides, prior methods often repurpose existing benchmarks (e.g., SCAN; Lake and Baroni 2018, ARC; Chollet 2019) without modification, where potential data contamination may undermine validity as models could use memorized solutions.
To address these gaps, we introduce a novel task of robust rule induction, which challenges models to identify underlying functions from input-output exemplars containing controlled noise injections. Our approach features three innovations: (1) an evaluation pipeline from data synthesis to automatic evaluation, requiring models to output rules explicitly and enabling direct validation through programmatic execution. (2) A multi-dimensional robustness assessment with conventional accuracy and instance-level metrics. (3) A sample-steered refinement (SRR) methodology that improves both the capability and stability of LLMs’ inductive reasoning through diverse sampling and iterative diagnostic refinement based on rule execution.
@@ 2.1 Problem Definitions
We formulate robust rule induction as the problem of identifying latent mapping functions from input-output pairs while maintaining consistent reasoning performance under noisy conditions. While prior works Wang et al. (2024); Liu et al. (2024) focus on basic rule discovery capabilities, we specifically investigate the robustness of inductive reasoning when the observations contain conflicting patterns. Some examples are shown in Figure 1. The observations may contain both normal and noisy examples, and the model is required to infer the correct rule regardless of the noise.
@@ 3 Sample-steered Rule Refinement
Motivated by previous methods Qiu et al. (2024); Wang et al. (2024) using iterative hypothesis refinement, we propose Sample-steered Rule Refinement (SRR), a framework to improve inductive reasoning in both clean and noisy conditions. This method conducts inductive reasoning through three-phase optimization: (1) contrastive hypothesis generation to bootstrap diverse rule candidates, (2) diagnostic feedback construction through evidence-aware sampling, and (3) iterative rule refinement using failure-driven corrections. The full algorithm is detailed in Algorithm 1.
@@
Different from previous methods based on the whole seen examples, we first generate hypotheses from random subsets of the seen examples, plus one hypothesis using the full observations. This step ensures coverage of both local patterns and global consistency. The number of examples in the subset is less than the full seen set, thus the noise in the subset has a greater impact. The sampled subset may also contain completely noise-free examples, which can better distinguish the rules.
@@
After generating the initial hypotheses, the induced rule is applied to the seen examples through code interpreter. We collect the correct and incorrect cases and sample these cases as feedback. We focus more on the incorrect cases, as they provide more information about the rule refinement, while the correct cases are used as positive feedback to reinforce the rule. The hypothesis with the highest accuracy on the seen examples is selected as the initial rule.
@@
At each iteration, the model receives the current selected rule with sampled feedback to generate a new rule. The new hypothesis is compared to the previous one, and the more accurate one is selected. This iterative process continues until the accuracy of the rule on the seen examples exceeds a predefined threshold or the maximum number of iterations is reached.
@@ 4.2 Datasets
We use three datasets with their subsets: Arithmetic, Cryptography, and List Functions. These datasets offer different rule induction challenges, including mathematical calculations, symbol representation, and list operations. To ensure that the noise reflects plausible perturbations encountered in real-world scenarios, we balance realism with experimental control while minimizing artificiality in designing the noise. Comprehensive description and statistics are shown in Table 1.
Dataset Noise Construction Subset # Instance Arith. Decimal Addition 7-base 8-base 9-base Crypto. Random Replacement Caesar Atbash Keyboard List Func. Random Replacement N/A
@@
The arithmetic task, introduced by Wu et al. (2024), involves counterfactual two-digit addition in non-decimal bases. To modulate difficulty and prioritize reasoning over memorization, we focus on the base-7, 8, 9 systems, and the noisy example is the common 10-based equations. The input is two two-digit numbers in the corresponding base, connected by a plus sign, and the output is the sum of the two numbers. To ensure the model can reason correctly, we guarantee that there must be a carry-over in the addition process so that the model must reason about the carry-over instead of simply adding the numbers.
@@
We use three types of substitution ciphers: Caesar, Atbash, and Keyboard. The Caesar cipher is a cipher that shifts the alphabet by a fixed number of positions. The Atbash cipher is a cipher that replaces each letter with the letter symmetrically opposite in the alphabet (e.g., A to Z). The Keyboard cipher replaces letters according to their positions in the alphabet with the corresponding positions on the keyboard (e.g., A to Q, B to W). The input is a word, and the output is the encrypted text. We randomly replace the characters in the output with other characters to generate noisy examples.
@@
The list functions dataset Rule (2020) evaluates the concept learning ability in the domain of cognitive science. The task is to induce a function that maps a list of numbers to another list of numbers. Each instance in the dataset corresponds to a list manipulation operation, such as sorting, reversing, or filtering. We generate the examples of each instance under controlled conditions to ensure the rule can be induced. The noisy examples are generated by randomly replacing the numbers in the output list with other numbers.
@@ 4.3 Robustness Under Different Noise Levels
First, contrary to conventional assumptions, noise introduction does not universally degrade performance, instead, the models exhibit performance fluctuations, demonstrating their inherent sensitivity to conflicting patterns. For example, GPT-4o achieves improved accuracy on Caesar cipher tasks at noise ( absolute improvement over clean data), while List Functions exhibit monotonic performance decay with increasing noise levels. In some cases, moderate noise even improves performance.
@@
Model Method Arithmetic Cryptography List Functions 7-base 8-base 9-base Caesar Atbash Keyboard GPT-4o DO CoT SC SR HR SRR Deepseek-V3 DO CoT SC SR HR SRR
Dataset GPT-4o DeepSeek-V3 SC HR-0 SRR-0 SC HR-0 SRR-0 Arithmetic Arithmetic Arithmetic Crypto Crypto Crypto List Functions Average
@@ 4.4 Method-wise Effectiveness Comparison
Direct Output may limit the model’s reasoning ability, as the model must output the rule directly without any intermediate steps. To assess the robustness of different reasoning paradigms, we compare our Sample-steered Rule Refinement (SRR) method with three reasoning methods: (1) Chain of Thought (CoT; Wei et al. 2022), which decomposes reasoning into step-by-step rationales; (2) Self-Consistency (SC; Wang et al. 2023), which aggregates multiple CoT trajectories through majority voting; (3) Self-Refine (SR; Madaan et al. 2023), which iteratively improves hypotheses using self-generated feedback; and (4) Iterative Hypothesis Refinement (HR; Qiu et al. 2024), which iteratively generates and ranks multiple hypotheses at each step, then refines the best hypothesis with feedback from incorrect predictions. Table 3 presents the task accuracy under noise and the deviation from clean data for GPT-4o and DeepSeek-V3.
@@
As shown in Table 3, our SRR framework achieves state-of-the-art performance across out of task-model combinations while exhibiting relatively low performance degradation ( average gap). This dual advantage stems from two mechanisms: (1) Diversity-aware hypothesis generation explores broader solution spaces through subset sampling, outperforming SC’s majority voting and the Hypothesis Refinement Qiu et al. (2024) that amplifies the similar patterns at the rule initialization stage (Table 4); (2) Execution-guided feedback leverages code interpreter for objective error detection, circumventing LLMs’ inherent deductive limitations Chen et al. (2023b); Cheng et al. (2024); Zheng et al. (2025a).
@@
While task accuracy provides comprehensive insights, consistency scores uncover fundamental reasoning fragility. As depicted in Figure 4, the Atbash cipher task exhibits particularly low consistency despite modest accuracy changes. The slight variation in accuracy may merely be an illusion created by the combined effects of noise interference (solvedunsolved cases) and incorrect generalization (unsolvedsolved cases). Prior and concurrent works’ singular focus on accuracy fluctuations Qiu et al. (2024); Zhou et al. (2024); Mirzadeh et al. (2024); Huang et al. (2025a) overlooks this critical duality in reasoning robustness.
@@
We observe dramatic performance gaps on counterfactual tasks like base-7 and base-9 arithmetic ( vs. for DeepSeek-V3 vs. GPT-4o) and Keyboard ciphers. Diagnostic analysis of responses from LLMs shows models default to familiar rules rather than induction. GPT-4o persistently misinterprets base-7 and base-9 addition as “base-8” or “decimal sum with constant”, while Atbash and Keyboard ciphers get erroneously classified as Caesar shifts in the reasoning process or self-generated feedback. Although Atbash and Caesar ciphers exhibit similar levels of complexity in transformation, their performance gaps are significant. These failures reveal that there is a pattern overfitting in the reasoning process. Current models’ inductive reasoning essentially operates through pattern matching rather than abstract induction. When the scarcity of counterfactual tasks in training data forces models to rely on “real” rule induction with unseen rules rather than recitation, model performance plummets dramatically.
@@ 4.5 Extended Explorations
Comparison with DeepSeek-R1
We focus on the Atbash cipher and List Functions. Atbash provides unified task semantics yet challenges models with unfamiliar transformation logic, while List Functions captures diverse rule abstraction scenarios. Table 5 compares task accuracy and consistency scores under clean and noise conditions. Notably, DeepSeek-R1 achieves higher task accuracy on both tasks than DeepSeek-V3. However, its consistency scores are still not competitive enough, indicating unresolved instability. Detailed breakdowns of consistency scores are shown in 6. For Atbash, the comparable number of right-to-wrong and wrong-to-right suggests instability and randomness in its reasoning process. Manual inspection shows DeepSeek-R1 still interprets Atbash cipher as character shifts in failed cases, mirroring previous pattern-overfitting behavior. For List Functions, the elevated right-to-wrong rate indicates its sensitivity to input perturbations.
Dataset Method Clean Acc Noise Acc Consistency Atbash CoT SRR List Func. CoT SRR
Dataset Method # BothR # BothW # RtoW # WtoR Atbash CoT SRR List Func. CoT SRR
@@
Rule (2020) reports human performance on List Functions. We divide the tasks into three difficulty levels in a ratio based on the sorted mean human performance. To contrast with human-like thinking patterns, we analyze consistency across 12 trials (4 noise levels 3 runs) in Section 4.3. Figure 5 visualizes the distribution of consistency and performance per task in List Functions, stratified by difficulty. LLMs show macro-level alignment yet micro-level divergence compared to human. While LLMs broadly mirror human-like stability, showing higher consistency on both simple and hard tasks, with moderate instability on medium-difficulty tasks, their internal patterns reveal critical deviations. For hard tasks, models display unpredictability (e.g., sporadic success) rather than systematic incapacity. Simple tasks, despite their low complexity, exhibit in consistent performance unrelated to intrinsic difficulty. Notably, GPT-4o and DeepSeek-V3 demonstrate similar behavioral trends, suggesting a shared inductive bias.
@@ 5 Discussion
LLMs heavily rely on prior knowledge rather than pure induction. They blend memorized pattern matching with shallow reasoning. When observations align with priors (e.g., Caesar cipher), they demonstrate pseudo-robustness. However, conflicting patterns and counterfactual scenarios expose this fragility—models either default to familiar templates or enter unstable hypothesis oscillations, as shown in Table 7. Even the state-of-the-art reasoning model exhibits unstable performance in unified tasks. This contrasts sharply with human few-shot learning, where true rule abstraction enables stable generalization Lake et al. (2015). Achieving human-level robust reasoning and few-shot induction requires disentangling knowledge recitation from inductive rule formation.
Unstable Hypothesis Oscillation |
Seen Examples (base-9 arithmetic): I: 68+68 O: 147; |
I: 80+42 O: 132; I: 76+17 O: 104; I: 57+56 O: 113; |
I: 38+14 O: 53; I: 43+84 O: 137; I: 48+55 O: 114; … |
Answer: To solve this problem, let’s analyze the examples to identify a potential pattern: 1. 68 + 68 = 147: The expected output of 68 + 68 should be 136, but the output is 147. This suggests there is a transformation applied to the arithmetic result. 2. 80 + 42 = 132: Normally, 80 + 42 equals 122, but the output is 132. … 4. 57 + 56 = 113: The expected result is 113, which matches the output. … The examples seem to suggest a pattern of adding an extra 10 to the arithmetic sum of the inputs, except in the case where the sum already matches the desired output (as in the case of 57 + 56 = 113). With these observations, we can propose a rule: If the sum of the numbers is not equal to the expected output, add 10 to the arithmetic sum. |
Default to Familiar Templates |
Seen Examples (Atbach cipher): I: cuissard O: xfrhhziw; |
I: tripsis", O: girkhrh; I: Maccabaeus O: Nzxxzyzvfh; … |
Answer: To solve this problem, let’s analyze the given examples to find a pattern: 1. cuissard->xfrhhziw 2. tripsis->girkhrh 3. Maccabaeus->Nzxxzyzvfh … Observations: 1. Character Shifts: The transformation seems to involve shifting characters. For example, ‘c’ in “cuissar” becomes ‘x’ in “xfrhhziw”. This suggests a possible Caesar cipher or similar character shift. … 4. Pattern Recognition: The shifts are not consistent across all examples, suggesting that some examples might be incorrect or that multiple rules are being applied. However, a common shift that appears frequently is 21 positions backward (e.g., ‘c’ to ’x’, ‘P’ to ‘K’, ‘C’ to ‘X’, ‘W’ to ‘D’). Hypothesis: Given the frequent appearance of a 21-position backward shift, let’s hypothesize that the rule involves shifting each character 21 positions backward in the alphabet. |